OpenStudy (itrymath):

Which of the following statements best describes the effect of replacing the graph of y = f(x) with the graph of y = f(x) - 9? (3 points) The graph of y = f(x) will shift up 9 units. The graph of y = f(x) will shift down 9 units. The graph of y = f(x) will shift left 9 units. The graph of y = f(x) will shift right 9 units.

OpenStudy (itrymath):

@ShadowLegendX

OpenStudy (itrymath):

C

OpenStudy (mww):

you are taking 9 away from every function value. So the whole graph is 9 less than before and will be vertical downshift of 9 units.

OpenStudy (itrymath):

so B ?

OpenStudy (itrymath):

if so im puzzled

OpenStudy (mww):

what's puzzling?

OpenStudy (itrymath):

i was so sure the answer was C

OpenStudy (mww):

The rules for transformations f(x) + k moves graph up k units f(x) - k moves graph down k units f(x-R) moves graph right R units f(x+L) moves graph left L units (k, L, R are all positive) Try drawing these to verify Draw y = x^2 and y = x^2 - 4 and see the difference.

OpenStudy (itrymath):

oh okay i see

OpenStudy (mww):

The difference between f(x) + k and f(x-k) is that for the former, you find the function at x then you add k units inducing a vertical shift. f(x-k) you take k units away from each x value then apply the function after inducing a horizontal shift

OpenStudy (itrymath):

okay so just to be clear answer is B

OpenStudy (itrymath):

@mww

OpenStudy (mww):

yes